The accounting department of a large limousine company is analyzing the costs of its services. The cost data and level of activity for the past 16 months follow.

Month Special Analyses Customer
Accounts
Paychecks
Processed
Accounting
Service Costs
1 3 300 1,240 $ 61,000
2 2 260 1,120 58,300
3 4 270 1,280 61,500
4 3 290 1,230 60,000
5 3 240 890 58,000
6 3 320 930 61,500
7 4 260 1,140 59,700
8 3 280 820 59,700
9 2 160 970 58,900
10 2 160 970 60,500
11 1 330 870 60,100
12 1 330 1,020 60,500
13 3 350 1,140 61,500
14 4 340 930 61,000
15 3 310 1,240 60,300
16 3 210 840 61,800
Totals 44 4,410 16,630 $ 964,300

In addition to the above information, you learn that the accounting department had the following total costs for the past 16 months for each of the following.

Total cost of paychecks processed $ 194,571
Total cost of maintaining customer accounts 134,946
Total cost of performing special analyses 173,360
Total fixed costs (total for 16 months) 461,423
Total costs $ 964,300

Required:

a-1. What is the cost per unit for paychecks processed?

a-2. What is the cost per unit for customer accounts maintained?

a-3. What is the cost per unit for special analyses performed?

b. Assuming the following level of cost-driver volumes for a month, what are the accounting department's estimated costs of doing business using the account analysis approach?

• 950 paychecks processed.

• 220 customer accounts maintained.

• 4 special analyses.
Answer #1
a-1. Cost per unit for paychecks processed=Total cost of paychecks processed/Total number of paychecks processed=194571/16630=$ 11.7
a-2. Cost per unit for customer accounts maintained=Total cost of maintaining customer accounts/Total number of customer accounts=134946/4410=$ 30.6
a-3. Cost per unit for special analyses performed=Total cost of performing special analyses/Total number of special analyses=173360/44=$ 39460
b. Estimated cost:
$
Cost of paychecks processed (950*11.7) 11115
Cost of maintaining customer accounts (220*30.6) 6732
Cost of performing special analyses (4*39460) 157840
Fixed cost per month (461423/16) 28839
Total 204526